OpenStudy (anonymous):

r2 - 14r + ____? help please

OpenStudy (anonymous):

Provide a little more information? The question doesn't quite make sense as is.

OpenStudy (anonymous):

it has to do with perfect square trinomial.

OpenStudy (anonymous):

(ax+b)^2=(a^2)(x^2) + 2bx + b^2 In this case a^2 = 1 and 2b = -14

OpenStudy (anonymous):

thank you!!